Answer parts a through e regarding the following investment project.

Ms. Phy Nance, a project analyst at DROF Motor Co., would like her project included in next year’s capital budget. According to her report, the new equipment would cost $185,217 and its projected cash flows would be as follows:

Year 140,000

Year 250,000

Year 353,000

Year 453,000

Year 570,000

The cost of capital is 12%

a) What is the NPV calculated by the analyst? If her numbers are correct should you accept the project?
b) Assume that the analyst included shipping and set-up costs of $5,217 in her estimation of the cost of the initial investment. Is this procedure correct? Why?
c) Assume that no consideration was given to the fact that the floor space used for this project could have been rented each year for $9,000 per year. How would this affect the calculation?
d) Included in her analysis, Ms. Nance calculated that lighting, heating and air conditioning cost the company $7,500 per year and that the new equipment occupies 40% of the total floor space --so she deducted $3,000 per year from the cash flows to reflect overhead costs for the new equipment. Is this deduction correct or should it be reversed?
e) Given your answers to b, c, and d, what is your calculation of Net Present Value for this project and would you include it in next year’s capital budget?
Answer #1

Below is the cash flow as given in the question:

Year 0 (-185,217) Year 1 - 40,000 Year 2 - 50,000 Year 3 - 53,000 Year 4 - 53,000 Year 5 - 70,000

a) NPV = -C0 + C1 / [ (1 + i)^1 ] +C2 / [ (1 + i)^2 ] +C3 / [ (1 + i)^3 ] +C4 / [ (1 + i)^4 ] +C5 / [ (1 + i)^5 ]

= -185,217+ 40,000 / [ (1 + 0.12)^1 ] +50,000 / [ (1 + 0.12)^2 ] +53,000 / [ (1 + 0.12)^3 ] +53,000 / [ (1 + 0.12)^4 ] +70,000 / [ (1 +0.12)^5 ]

= -185,217 + 35714.28 + 39,859.69 + 37724.35 + 33682.45 + 39719.87

= 1483.64

We can also do this NPV calculation in excel easily. Since NPV is positive we should accept the project

b. Yes this procedure is correct. These are the cost needed to start the project and are classified as initial investment outlay. Therefore, these are relevant cashflows for the analysis.

C. This is the opportunity cost and should be included. The cash flow for each year will be subtracted by $9,000

Year 0 = - 185,217

Year 1 - 40,000 - $9,000 = 31,000

Year 2 - 50,000 - $9,000 = 41,000

Year 3 - 53,000 - $9,000 = 44,000

Year 4 - 53,000 - $9,000 = 44,000

Year 5 - 70,000 - $9,000 = 61,000

Now, Calculate NPV with the new cash flow and with above formula as used for question "a" or Excel, we will get NPV as,

NPV = -C0 + C1 / [ (1 + i)^1 ] +C2 / [ (1 + i)^2 ] +C3 / [ (1 + i)^3 ] +C4 / [ (1 + i)^4 ] +C5 / [ (1 + i)^5 ]

= -185,217+ 31,000 / [ (1 + 0.12)^1 ] +41,000 / [ (1 + 0.12)^2 ] +44,000 / [ (1 + 0.12)^3 ] +44,000 / [ (1 + 0.12)^4 ] +61,000 / [ (1 +0.12)^5 ]

= -185,217 + 27678.57 + 32684.94 + 31318.33 + 27962.79+ 34613.03

= -30,959.34

Still NPV is Negative and we should not accept the project.

d) This is not correct and should be revered. All overhead costs are not relevant in NPV analysis. Only the incremental overhead costs due to a project are relevant. Overhead costs are not relevant if total overhead cash flows remain the same

e) After adjustment below will be the cashflows

Year 0 = - 185,217

Year 1 - 40,000 - $9,000 + 3000= 34,000

Year 2 - 50,000 - $9,000 + 3000= 44,000

Year 3 - 53,000 - $9,000 +3,000= 47,000

Year 4 - 53,000 - $9,000 +3,000= 47,000

Year 5 - 70,000 - $9,000 + 3,000 = 64,000

NPV = -C0 + C1 / [ (1 + i)^1 ] +C2 / [ (1 + i)^2 ] +C3 / [ (1 + i)^3 ] +C4 / [ (1 + i)^4 ] +C5 / [ (1 + i)^5 ]

= -185,217+ 34,000 / [ (1 + 0.12)^1 ] +44,000 / [ (1 + 0.12)^2 ] +47,000 / [ (1 + 0.12)^3 ] +47,000 / [ (1 + 0.12)^4 ] +64,000 / [ (1 +0.12)^5 ]

= -185,217+ 30357.14 + 35076.53+ 33453.67+ 29869.34+ 36315.31

= -20,145

Therefore, project should not be accepted